# Solving adversarial examples requires solving exponential misalignment

### 03 Mar 2026 v1

@@ Abstract
Figure 1: Visualization of our main argument: We show that a network’s perceptual manifold (PM) for any class concept (e.g. cat), defined to be the set of all images confidently perceived by the network as that class, is extremely high dimensional: $3000$ out of a total of $3072$ for CIFAR10 (large red manifold) and $\approx 135,000$ out of $150,528$ for CLIP and ImageNet. In contrast natural images perceived by humans as any class (e.g. dogs, airplanes or cats (blue, green and bright red manifolds)) are only $\approx 20$ dimensional. This indicates that machine and human PMs for any concept are exponentially misaligned: there are exponentially many inputs confidently perceived as any given concept by machines, but not by humans (e.g. the two noise images in the network’s cat perceptual manifold). This exponential misalignment also explains the origin of adversarial examples: e.g. because the network’s cat PM fills up so much of image space, any other input (e.g. dog or airplane) is extremely close to it.
@@ 2.1 A network’s perceptual manifold for a class concept
where $p_{0}$ is the threshold confidence level. Unless otherwise stated, we will choose a confidence level of $p_{0}=0.9$, i.e. roughly the average clean accuracy of the various models on CIFAR10 (Krizhevsky, ) test images). Crucially, the perceptual manifold characterizes the subset of the input space that the network confidently perceives as exemplars of concept $c$. A main goal is to explore the misalignment between this machine PM and the human PM of natural images that humans would confidently classify as exemplars of concept $c$.
@@ 3.1 Non-robust networks are exponentially misaligned
To show that standard non-robust networks are exponentially misaligned, in Fig. 2 we compare the dimensionality of the set of natural images to that of the PM of a standard non-robust model for each class of CIFAR-10. The latter exceeds the former by orders of magnitude: the Participation Ratio (PR) and Two Nearest Neighbors (2NN) estimates for natural images are $\approx 10$ and $\approx 20$, while those of the model’s PM are $\approx 3000$ and $\geq 500$ respectively (noting that the latter is a lower bound). Again, since volume typically grows exponentially with dimension, this orders of magnitude dimensional discrepancy between machine and natural image PMs indicates exponential misalignment: there are exponentially many images artificial networks confidently assign to any concept that humans do not (see e.g. Fig. 8). We repeat this analysis on ImageNet-1K in Section H.1 with consistent results. Crucially, this catastrophic dimensional gap is not an artifact of low-resolution data; on ImageNet-1K, the PM of a standard ResNet-50 occupies over 130,000 out of 150,528 dimensions, while the natural image concepts remain roughly 20-dimensional.
@@ 3.2 Geometry relating misalignment to fragility
We now provide a logical connection between a networks PM dimensionality and the network’s fragility to bounded norm adversarial attacks. Intuitively, the higher the dimension of a manifold, the more of ambient space it will fill up, and the closer any point in the ambient space will be to that manifold. In the context of CIFAR10 images, if the network’s PM is $3000$ dimensional in an ambient space of $3072$ dimensions, it fills up nearly all of image space, and so any image $x\in[0,1]^{D}$ not in the PM, should be very close to the PM, and therefore is highly likely to admit a successful bounded norm adversarial attack.
@@ Appendix H ImageNet Experiments
To verify that the results reported in the main text are not an artifact of low-resolution datasets like CIFAR-10, we replicated our core analyses on ImageNet-1K (Russakovsky et al., 2015). We once again utilized pretrained models from the RobustBench leaderboard (Croce et al., 2021), restricting our analysis to models with a clean accuracy of $\geq 70\%$. This constraint isolates the geometric effects of robustness from general baseline (clean) performance differences. Due to the severe computational constraints of high-dimensional optimization, we exclusively use the Participation Ratio (PR) to estimate dimensionality and set our target confidence threshold to $p_{0}=0.8$, reflecting the average clean accuracy of this model cohort.

### 11 Mar 2026 v2

@@ Abstract
Figure 1: Visualization of our main argument: We show that a network’s perceptual manifold (PM) for any class concept (e.g. cat), defined to be the set of all images confidently perceived by the network as that class, is extremely high dimensional: $3060$ out of a total of $3072$ for CIFAR10 (large red manifold) and $\approx 135,000$ out of $150,528$ for CLIP and ImageNet. In contrast natural images perceived by humans as any class (e.g. dogs, airplanes or cats (blue, green and bright red manifolds)) are only $\approx 20$ dimensional. This indicates that machine and human PMs for any concept are exponentially misaligned: there are exponentially many inputs confidently perceived as any given concept by machines, but not by humans (e.g. the two noise images in the network’s cat perceptual manifold). This exponential misalignment also explains the origin of adversarial examples: e.g. because the network’s cat PM fills up so much of image space, any other input (e.g. dog or airplane) is extremely close to it.
@@ 2.1 A network’s perceptual manifold for a class concept
where $p_{0}$ is the threshold confidence level. Unless otherwise stated, we will choose a confidence level of $p_{0}=0.9$, which is much higher than that of robust networks ($60\%$ - $70\%$ on correctly classified CIFAR10 (Krizhevsky, ) test images) and comparable to that of standard non-robust networks (often $\geq 95\%$). Crucially, the perceptual manifold characterizes the subset of the input space that the network confidently perceives as exemplars of concept $c$. A main goal is to explore the misalignment between this machine PM and the human PM of natural images that humans would confidently classify as exemplars of concept $c$.
@@ 3.1 Non-robust networks are exponentially misaligned
To show that standard non-robust networks are exponentially misaligned, in Fig. 2 we compare the dimensionality of the set of natural images to that of the PM of a standard non-robust model for each class of CIFAR-10. The latter exceeds the former by orders of magnitude: the Participation Ratio (PR) and Two Nearest Neighbors (2NN) estimates for natural images are $\approx 10$ and $\approx 20$, while those of the model’s PM are $\approx 3060$ and $\geq 500$ respectively (noting that the latter is a lower bound). Again, since volume typically grows exponentially with dimension, this orders of magnitude dimensional discrepancy between machine and natural image PMs indicates exponential misalignment: there are exponentially many images artificial networks confidently assign to any concept that humans do not (see e.g. Fig. 8). We repeat this analysis on ImageNet-1K in Section H.1 with consistent results. Crucially, this catastrophic dimensional gap is not an artifact of low-resolution data; on ImageNet-1K, the PM of a standard ResNet-50 occupies over 130,000 out of 150,528 dimensions, while the natural image concepts remain roughly 20-dimensional.
@@ 3.2 Geometry relating misalignment to fragility
We now provide a logical connection between a networks PM dimensionality and the network’s fragility to bounded norm adversarial attacks. Intuitively, the higher the dimension of a manifold, the more of ambient space it will fill up, and the closer any point in the ambient space will be to that manifold. In the context of CIFAR10 images, if the network’s PM is $3060$ dimensional in an ambient space of $3072$ dimensions, it fills up nearly all of image space, and so any image $x\in[0,1]^{D}$ not in the PM, should be very close to the PM, and therefore is highly likely to admit a successful bounded norm adversarial attack.
@@ Appendix H ImageNet Experiments
To verify that the results reported in the main text are not an artifact of low-resolution datasets like CIFAR-10, we replicated our core analyses on ImageNet-1K (Russakovsky et al., 2015). We once again utilized pretrained models from the RobustBench leaderboard (Croce et al., 2021), restricting our analysis to models with a clean accuracy of $\geq 70\%$. This constraint isolates the geometric effects of robustness from general baseline (clean) performance differences. Due to the severe computational constraints of high-dimensional optimization, we exclusively use the Participation Ratio (PR) to estimate dimensionality and set our target confidence threshold to $p_{0}=0.8$, once again higher than that of robust networks ($60\%-70\%$ on correctly classified images) and comparable to that of standard non robust networks (often $\geq 85\%$).